# On sporadic symmetry breaking operators for principal series representations of the de Sitter and Lorentz groups

### v1

@@ 3.3 Reduction Theorems
For any $\mu\in{\mathbb{C}}$ and any $\ell\in{\mathbb{N}}$, the imaginary Gegenbauer differential operator $S_{\ell}^{\mu}$ is defined as follows (see (8.8)) (cf. [20, Eq. (6.5)]):
@@
Now, for $f_{0},f_{\pm 1},f_{\pm 2}\ldots,f_{\pm N}\in{\mathbb{C}}[t]$ set $\mathbf{f}:={}^{t}(f_{-N},\ldots,f_{0},\ldots,f_{N})$ and let $L_{j}^{A,\pm}(\mathbf{f})(t)$ and $L_{j}^{B,\pm}(\mathbf{f})(t)$ $(j=0,\dots,N)$ be the following polynomials:
$\begin{array}[]{l}\begin{rcases*}L_{j}^{A,+}(\mathbf{f})(t):=S_{a+m-j}^{\lambda+j-1}f_{j}-2(N-j)\frac{d}{dt}f_{j+1}.\\[6.0pt] L_{j}^{A,-}(\mathbf{f})(t):=S_{a-m-j}^{\lambda+j-1}f_{-j}+2(N-j)\frac{d}{dt}f_{-j-1}.\end{rcases*}(j=0,1,\ldots,N)\\[30.0pt] \begin{rcases*}L_{j}^{B,+}(\mathbf{f})(t):=2(-m(\lambda+a-1)+j(\lambda-1+\vartheta_{t}))f_{j}\\ +(N-j)\frac{d}{dt}f_{j+1}+(N+j)\frac{d}{dt}f_{j-1}.\\[6.0pt] L_{j}^{B,-}(\mathbf{f})(t):=2(m(\lambda+a-1)+j(\lambda-1+\vartheta_{t}))f_{-j}\\ -(N+j)\frac{d}{dt}f_{-j+1}-(N-j)\frac{d}{dt}f_{-j-1}.\end{rcases*}(j=1,\ldots,N)\end{array}$ (3.15)
@@
Given $\mathbf{g}=(g_{k_{d}})_{d=0}^{2N}={}^{t}(g_{k_{0}},\ldots,g_{k_{2N}})$ with $g_{k_{d}}\in{\mathbb{C}}[t]$, we define the “reflection”  $\mathbf{g}^{\text{ref}}$ of $\mathbf{g}$ as follows:
$\mathbf{g}^{\text{ref}}={}^{t}(g_{-N}^{\text{ref}},\ldots,g_{N}^{\text{ref}}):=\begin{pmatrix}&&1\\ &\iddots&\\ 1&&\end{pmatrix}\mathbf{g}={}^{t}(g_{k_{2N}},\ldots,g_{k_{0}}).$
@@
The following three conditions on the quadruple $(\lambda,a,N,m)\in{\mathbb{C}}\times{\mathbb{N}}_{\geq N-m}\times{\mathbb{N}}\times{\mathbb{N}}_{>N}$ are equivalent.
@@
In this section we prove Theorem 3.6, which is the key result to show both Theorems 1.3 and 1.4. In other words, we solve the system (3.17), which is given by the following differential equations:
@@
$f_{\pm j}\in\text{{Pol}}_{a-m\pm j}[t]_{\text{{{even}}}}\subseteq{\mathbb{C}}[t]\enspace(j=0,1,\ldots,N).$ (4.2)
Even though this system is formed by linear ordinary differential equations of order at most 2 with polynomial solutions, it is an overdetermined system, since there are $2(2N+1)$ equations and $12N+$ 1 functions. This makes it difficult to solve, specially since there is not a clear relation between equations of type A and B.
@@
    Inductively, for $j=0,\dots,N-1$, obtain $f_{\pm j}$ by solving $(B_{j}^{\pm})$. After that, check that these $f_{\pm j}$ satisfy equations $(A_{j}^{\pm})$. For $f_{0}$ we obtain two different expressions (one coming from the $+$ side and another coming from the -- side).
@@
The three-phase strategy explained above was also used to solve the system (4.1) when $m=N$ (see [36]). However, when $m>N$, the situation changes considerably, and even if the same idea works, the procedure is more complicated; specially Phase 3. One aspect that illustrates the fundamental difference between the cases $m=N$ and $m>N$, is that while in the former case the system (4.1) has non-zero solution for any $(\lambda,\nu)\in{\mathbb{C}}^{2}$ satisfying $\nu-\lambda\in{\mathbb{N}}$, when $m>N$ non-zero solutions do not exist if $\lambda\notin{\mathbb{Z}}$ (see Proposition 4.1 below).
@@ 4.1 Phase 1: Obtaining $f_{\pm N}$ up to constant–Necessary condition on $\lambda$
Since the system (4.1) does not have non-zero solutions for $a<m-N$ (see Proposition 3.3), from now on we suppose that $a\geq m-N$. Recall that we are also under the assumption $m>N$, so that $a\in{\mathbb{N}}_{+}$.
@@
$(A_{N}^{+})\enspace S_{a+m-N}^{\lambda+N-1}f_{N}=0,$ $(A_{N}^{-})\enspace S_{a-m-N}^{\lambda+N-1}f_{-N}=0.$
@@
So a priori the degree of the polynomial $\widetilde{C}_{a+m-N}^{\lambda+N-1}(it)$ is greater than the degree of $f_{N}$. We will see that $\widetilde{C}_{a+m-N}^{\lambda+N-1}(it)$ actually has degree $a-m+N$ for an appropriate parameter $\lambda\in{\mathbb{C}}$ (in other words, that the higher terms of $\widetilde{C}_{a+m-N}^{\lambda+N-1}(it)$ vanish), so that the expression for $f_{N}$ is indeed correct. If we take a closer look at the definition of the renormalized Gegenbauer polynomial (8.7), we deduce that this can only happen if $\lambda\in{\mathbb{Z}}$ (an example of this type of phenomenon can be seen in Lemmas 8.8 and 8.10). The next result shows that, as a matter of fact, the system does not have non-zero solutions if $\lambda\notin{\mathbb{Z}}$.
@@
for some $q_{j}^{+}\in{\mathbb{C}}$. By the condition (4.2), we have that $f_{j}\in\text{{Pol}}_{a-m+j}[t]_{\text{even}}$. However, since $m>N$, we have $a-m+j<a+m-j$. That is,
@@
Thus, $f_{-j}$ is constant. By (4.2) we deduce that if $a-m-j$ is odd, $f_{-j}=0$ necessarily. Suppose then that $a-m-j$ is even and set $f_{-j}=q_{j}^{-}\in{\mathbb{C}}$. By $(A_{j}^{-})$ we have that
@@
00 0 $=4(m(\lambda+a-1)+j(\lambda-1+\vartheta_{t}))f_{-j}-2(N+j)\frac{d}{dt}f_{-j+1}-2(N-j)\frac{d}{dt}f_{-j-1}$ (4.4)
@@
Since $a-m-j$ is even, and by assumption we have
@@
As before, since $m>N$ we have $a-m+N<a+m-N$. Thus, from (4.3) and from Lemma 8.8, we deduce that $f_{N}=0$ necessarily. Now, by using Lemma 4.2(1) repeatedly, we have
@@ 4.2 Phase 2: Obtaining all $f_{\pm j}$ up to constant
$\Gamma_{j}^{+}$ $:=\frac{\Gamma\left(\lambda+N-1+\left[\frac{a+m-N+1}{2}\right]\right)}{\Gamma\left(\lambda+N-1+\left[\frac{a+m-2N+j+1}{2}\right]\right)}=\frac{\Gamma\left(\lambda+\left[\frac{a+m+N-1}{2}\right]\right)}{\Gamma\left(\lambda\left[\frac{a+m+j-1}{2}\right]\right)},$ (4.6)
$\Gamma_{j}^{-}$ $:=\frac{\Gamma\left(\lambda+N-1+\left[\frac{a-m-N+1}{2}\right]\right)}{\Gamma\left(\lambda+N-1+\left[\frac{a-m-2N+j+1}{2}\right]\right)}=\frac{\Gamma\left(\lambda+\left[\frac{a-m+N-1}{2}\right]\right)}{\Gamma\left(\lambda+\left[\frac{a-m+j-1}{2}\right]\right)}.$
@@
$\frac{1}{2}(-i)^{N-j}\frac{\Gamma(2N+1)}{\Gamma(N+j+1)}\Gamma_{j}^{+}=$ $\enspace\frac{1}{2}(-i)^{N-j-1}\frac{\Gamma(2N+1)}{\Gamma(N+j+2)}\Gamma_{j+1}^{+}$ (4.11)
×$\times(-i)(N+j+1)\gamma(\lambda+N-1,a+m-2N+j)\neq 0.$
@@
×$\times(N+j+1)C_{j+1,r}^{+}+((j-m)(\lambda+a-j-1)+2j(r+j-N))C_{j,r}^{+}\Big{]}\widetilde{C}_{a+m-2N+j+2r}^{\lambda+N-r-1}(it)$
@@
$-\frac{1}{2}(-i)^{N-j}\frac{\Gamma(2N+1)}{\Gamma(N+j)}\Gamma_{j}^{+}\frac{d}{dt}f_{j-1}=q_{N}^{+}\sum_{r=0}^{N-j+1}C_{j-1,r}^{+}\widetilde{C}_{a+m-2N+j+2r-2}^{\lambda+N-r}(it).$
@@
=\;= $2q_{N}^{+}\sum_{r=0}^{N-j}(N-j+1-r)C_{j-1,r}^{+}\Bigg{[}(\lambda+a+m+j-2+r-N)\widetilde{C}_{a+m-2N+j-1+2r}^{\lambda+N-1-r}(it)$
@@
We note that if $a+m+j-1$ is odd, then $c_{j-1}^{+}=0$ necessarily by (4.2). In particular, (4.14) above is satisfied. On the contrary, if $a+m+j-1$ is even, we have
@@
$\gamma(\lambda+N-1,a+m-N-s)=0,\enspace\text{ for some }s=1,2,\ldots,N.$
@@
$\gamma(\lambda+N-1,a-m-N-s)=0,\enspace\text{ for some }s=1,2,\ldots,N.$
@@
Indeed, if $a+m-j-2$ is odd, the left-hand side is equal to 1. And if $a+m-j-2$ is even, we have
$2\gamma(\lambda,a+m-j-2)=2\lambda+a+m-j-2=2m-j-N+s>s\geq 1$
@@ 4.3 Phase 3: Proving $f_{-0}=f_{+0}$
Let $(\lambda,N,a,m)\in{\mathbb{Z}}\times{\mathbb{N}}\times{\mathbb{N}}\times{\mathbb{Z}}$ with $m>N$, and suppose further that $a\geq m$. Then, the identity (4.19) holds if and only if the tuple $(\lambda,N,a,m,\widetilde{q_{N}^{\pm}})$ satisfies (4.20) and (4.21) below.
@@
Let $\alpha\equiv\alpha(\lambda,N,a,m)$ and $\beta\equiv\beta(\lambda,N,a,m)$ be the following polynomials on $\lambda$ of degree $N$
@@
$P(\lambda):=$ $\left(\lambda+\left[\frac{a+m-1}{2}\right]\right)\left[\frac{a+m}{2}\right]\alpha(\lambda,N,a,m)\beta(\lambda,N,a,-m)$ (4.24)
@@
The first one is straightforward, since $\alpha$ and $\beta$ have degree $N$ by (4.22) and (4.23), and since $Q$ is the product of $12N+$ 1 monomials.
@@
()($\Rightarrow$)( ⇒ ) Suppose that (4.19) holds. Let us show that both (4.20) and (4.21) hold. We divide this direction of the proof into the following three cases:
@@
$\begin{split}\delta_{\varepsilon+2}\fallingdotseq&\;(-1)^{m}\left(\lambda+\left[\frac{a+m-1}{2}\right]\right)\alpha(\lambda,N,a,m)\widetilde{q_{N}^{+}}\\ &-\left(\lambda+\left[\frac{a-m-1}{2}\right]\right)\frac{\Gamma(\left[\frac{a-m}{2}\right]+m)}{\Gamma(\left[\frac{a-m}{2}\right])}\alpha(\lambda,N,a,-m)\widetilde{q_{N}^{-}},\end{split}$ (4.31)
@@
The necessary and sufficient condition for this homogeneous system to have a non-zero solution $(\widetilde{q_{N}^{+}},\widetilde{q_{N}^{-}})$, is that the determinant of the matrix defining the system must be zero. In other words,
@@
×$\times{}_{3}F_{2}\left(\begin{matrix}-q,m-N,-N\\ -\left[\frac{a+m+2}{2}\right]+m-q+1\end{matrix},-2N;1\right).$
@@
(2) Now, suppose $m+N\leq a\leq m+2N+1$. If (4.19) is satisfied, then clearly $\delta_{a-m+2}=0$. Since the left-hand side of (4.19) has degree $a-m$, this term comes only from the right-hand side. A straightforward computation yields to
@@
×$\times\sum_{r=0}^{N}\begin{pmatrix}N\\ r\end{pmatrix}\frac{\Gamma(2N-r+1)}{\Gamma(N+1)}\frac{\Gamma(\lambda+a-N+r-1)}{\Gamma(\lambda+a-N-1)}.$
@@
[…]
@@
$1\displaystyle 1\leq 1+\left[\frac{N}{2}\right]\leq\left[\frac{a-m+2}{2}\right]\leq\left[\frac{2N+3}{2}\right]=N+$ 1 .
@@
Note that if $10<\eta<$ 1, $\beta_{\eta}(\lambda,N,a,\pm m)$ can be expressed by using the function ${}_{3}F_{2}$ as:
@@
()($\Leftarrow$)( ⇐ ) Let us show now that identity (4.19) is satisfied assuming that (4.20) and (4.21) hold. We do it in two steps. First, we make use of Lemma 8.10 and rewrite the right-hand side of (4.19). Then, we simplify both sides by using properties of the hypergeometric function ${}_{3}F_{2}$ to show that they actually coincide. From (4.21), we have that (4.19) amounts to:
@@
For $\lambda=N+1-a-q\in\Lambda(N,a)$ and $r=0,\ldots,N$, we set $b:=-\lambda-N+1+r=a+r+q-2N$ and $\ell:=a+m-2N+2r$. From $a\geq m$ and $q\geq\max(0,N-a+m)$, we clearly have $b,\ell\in{\mathbb{N}}$. Now, if $2b-\ell=a-m-2N+2q\geq 0$, then the right-hand side of (4.34) amounts to (4.35) due to Lemma 8.10. On the other hand, if $02b-\ell=a-m-2N+2q<$ 0, then (4.35) vanishes since all Gegenbauer polynomials in the sum have negative degrees. Similarly, the right-hand side of (4.34) is also zero. In fact, as $m\leq a<m+2N-2q$, we have
@@
We start by showing (4.37). By changing the order of the indexes in the sum $S_{1}$, we have
@@
[…]
@@
If $10\leq q\leq N-$ 1, (4.38) is trivially satisfied since both sides are zero. Suppose then $N\leq q\leq 2N$. If $d\geq N+1$, then $d-N-1\geq 0$. Therefore, $(-N)_{d}=(-N)(1-N)\cdots(d-N-1)=0$. In other words, the sum $S_{2}$ only has one term, namely, the one for $d=N$. Thus,

### v2

@@ 3.4 Reduction Theorems
For any $\mu\in{\mathbb{C}}$ and any $\ell\in{\mathbb{N}}$, the imaginary Gegenbauer differential operator $S_{\ell}^{\mu}$ is defined as follows (see (8.8)) (cf. [23, Eq. (6.5)]):
@@
[…]
$\begin{array}[]{l}\begin{rcases*}L_{j}^{A,+}(\mathbf{f})(t):=S_{a+m-j}^{\lambda+j-1}f_{j}-2(N-j)\frac{d}{dt}f_{j+1}.\\[6.0pt] L_{j}^{A,-}(\mathbf{f})(t):=S_{a-m-j}^{\lambda+j-1}f_{-j}+2(N-j)\frac{d}{dt}f_{-j-1}.\end{rcases*}(j=0,1,\ldots,N)\\[30.0pt] \begin{rcases*}L_{j}^{B,+}(\mathbf{f})(t):=2(-m(\lambda+a-1)+j(\lambda-1+\vartheta_{t}))f_{j}\\ \hskip 0.0pt plus 1.0+(N-j)\frac{d}{dt}f_{j+1}+(N+j)\frac{d}{dt}f_{j-1}.\\[6.0pt] L_{j}^{B,-}(\mathbf{f})(t):=2(m(\lambda+a-1)+j(\lambda-1+\vartheta_{t}))f_{-j}\\ \hskip 0.0pt plus 1.0-(N+j)\frac{d}{dt}f_{-j+1}-(N-j)\frac{d}{dt}f_{-j-1}.\end{rcases*}(j=1,\ldots,N)\end{array}$ (3.20)
@@
[…]
@@
Let $\lambda\in{\mathbb{C}}$, $a,N\in{\mathbb{N}}$ and $m\in{\mathbb{Z}}$ and suppose $a\geq N-m$ and $m>N$. Then, the following three conditions on the quadruple $(\lambda,a,N,m)$ are equivalent.
@@
In this section we prove Theorem 3.7, which is the key result to show both Theorems 1.3 and 1.5. In other words, we solve the system (3.22), which is given by the following differential equations:
@@
$f_{\pm j}\in\text{{Pol}}_{a-m\pm j}[t]_{\text{{{even}}}}\subseteq{\mathbb{C}}[t]\kern 4.62497pt(j=0,1,\ldots,N).$ (4.2)
Even though this system is formed by linear ordinary differential equations of order at most 2 with polynomial solutions, it is an overdetermined system, since there are $2(2N+1)$ equations and $2N+1$ functions. This makes it difficult to solve, especially since there is not a clear relation between equations of type A and B.
@@
    Inductively, for $j=0,\dots,N-1$, obtain $f_{\pm j}$ by solving $(B_{j}^{\pm})$. After that, check that these $f_{\pm j}$ satisfy equations $(A_{j}^{\pm})$. For $f_{0}$ we obtain two different expressions (one coming from the ++ side and another coming from the - side).
@@
The three-phase strategy explained above was also used to solve the system (4.1) when $m=N$ (see [39]). However, when $m>N$, the situation changes considerably, and even if the same idea works, the procedure is more complicated; especially Phase 3. One aspect that illustrates the fundamental difference between the cases $m=N$ and $m>N$, is that while in the former case the system (4.1) has non-zero solutions for any $(\lambda,\nu)\in{\mathbb{C}}^{2}$ satisfying $\nu-\lambda\in{\mathbb{N}}$, when $m>N$ non-zero solutions do not exist if $\lambda\notin{\mathbb{Z}}$ (see Proposition 4.1 below).
@@ 4.1 Phase 1: Obtaining $f_{\pm N}$ up to constant–Necessary condition on $\lambda$
Since the system (4.1) does not have non-zero solutions for a<mNa<m-N (see Proposition 3.3), from now on we suppose that $a\geq m-N$. Recall that we are also under the assumption $m>N$, so that $a\in{\mathbb{N}}_{+}$.
@@
$(A_{N}^{+})\kern 4.62497ptS_{a+m-N}^{\lambda+N-1}f_{N}=0,$ $(A_{N}^{-})\kern 4.62497ptS_{a-m-N}^{\lambda+N-1}f_{-N}=0.$
@@
So a priori the degree of the polynomial $\widetilde{C}_{a+m-N}^{\lambda+N-1}(it)$ is greater than the degree of $f_{N}$. We will see that $\widetilde{C}_{a+m-N}^{\lambda+N-1}(it)$ actually has degree am+Na-m+N for an appropriate parameter $\lambda\in{\mathbb{C}}$ (in other words, that the higher terms of $\widetilde{C}_{a+m-N}^{\lambda+N-1}(it)$ vanish), so that the expression for $f_{N}$ is indeed correct. If we take a closer look at the definition of the renormalized Gegenbauer polynomial (8.7), we deduce that this can only happen if $\lambda\in{\mathbb{Z}}$ (an example of this type of phenomenon can be seen in Lemmas 8.8 and 8.10). The next result shows that, as a matter of fact, the system does not have non-zero solutions if $\lambda\notin{\mathbb{Z}}$.
@@
for some $q_{j}^{+}\in{\mathbb{C}}$. By the condition (4.2), we have that $f_{j}\in\text{{Pol}}_{a-m+j}[t]_{\text{even}}$. However, since $m>N$, we have am+j<a+mja-m+j<a+m-j. That is,
@@
Thus, $f_{-j}$ is constant. By (4.2) we deduce that if amja-m-j is odd, $f_{-j}=0$ necessarily. Suppose then that amja-m-j is even and set $f_{-j}=q_{j}^{-}\in{\mathbb{C}}$. By $(A_{j}^{-})$ we have that
@@
$0$ $=4(m(\lambda+a-1)+j(\lambda-1+\vartheta_{t}))f_{-j}-2(N+j)\frac{d}{dt}f_{-j+1}-2(N-j)\frac{d}{dt}f_{-j-1}$ (4.4)
@@
Since amja-m-j is even, and by assumption we have
@@
As before, since $m>N$ we have am+N<a+mNa-m+N<a+m-N. Thus, from (4.3) and from Lemma 8.8, we deduce that $f_{N}=0$ necessarily. Now, by using Lemma 4.2(1) repeatedly, we have
@@ 4.2 Phase 2: Obtaining all $f_{\pm j}$ up to constant
$\Gamma_{j}^{+}$ $=\frac{\Gamma\left(\lambda+N-1+\left[\frac{a+m-N+1}{2}\right]\right)}{\Gamma\left(\lambda+N-1+\left[\frac{a+m-2N+j+1}{2}\right]\right)}=\frac{\Gamma\left(\lambda+\left[\frac{a+m+N-1}{2}\right]\right)}{\Gamma\left(\lambda\left[\frac{a+m+j-1}{2}\right]\right)},$ (4.6)
$\Gamma_{j}^{-}$ $=\frac{\Gamma\left(\lambda+N-1+\left[\frac{a-m-N+1}{2}\right]\right)}{\Gamma\left(\lambda+N-1+\left[\frac{a-m-2N+j+1}{2}\right]\right)}=\frac{\Gamma\left(\lambda+\left[\frac{a-m+N-1}{2}\right]\right)}{\Gamma\left(\lambda+\left[\frac{a-m+j-1}{2}\right]\right)}.$
@@
$\frac{1}{2}(-i)^{N-j}\frac{\Gamma(2N+1)}{\Gamma(N+j+1)}\Gamma_{j}^{+}=$ $\kern 4.62497pt\frac{1}{2}(-i)^{N-j-1}\frac{\Gamma(2N+1)}{\Gamma(N+j+2)}\Gamma_{j+1}^{+}$ (4.11)
$\times(-i)(N+j+1)\gamma(\lambda+N-1,a+m-2N+j)\neq 0.$
@@
$\times(N+j+1)C_{j+1,r}^{+}+((j-m)(\lambda+a-j-1)+2j(r+j-N))C_{j,r}^{+}\Big]\widetilde{C}_{a+m-2N+j+2r}^{\lambda+N-r-1}(it)$
@@
$\frac{1}{2}(-i)^{N-j}\frac{\Gamma(2N+1)}{\Gamma(N+j)}\Gamma_{j}^{+}\frac{d}{dt}f_{j-1}=q_{N}^{+}\sum_{r=0}^{N-j+1}C_{j-1,r}^{+}\widetilde{C}_{a+m-2N+j+2r-2}^{\lambda+N-r}(it).$
@@
$=\;$ $2q_{N}^{+}\sum_{r=0}^{N-j}(N-j+1-r)C_{j-1,r}^{+}\Bigg[(\lambda+a+m+j-2+r-N)\widetilde{C}_{a+m-2N+j-1+2r}^{\lambda+N-1-r}(it)$
@@
We note that if a+m+j1a+m+j-1 is odd, then $c_{j-1}^{+}=0$ necessarily by (4.2). In particular, (4.14) above is satisfied. On the contrary, if a+m+j1a+m+j-1 is even, we have
@@
$\gamma(\lambda+N-1,a+m-N-s)=0,\kern 4.62497pt\text{ for some }s=1,2,\ldots,N.$
@@
$\gamma(\lambda+N-1,a-m-N-s)=0,\kern 4.62497pt\text{ for some }s=1,2,\ldots,N.$
@@
Indeed, if a+mj2a+m-j-2 is odd, the left-hand side is equal to 1. And if a+mj2a+m-j-2 is even, we have
$2\gamma(\lambda,a+m-j-2)=2\lambda+a+m-j-2=2m-j-N+s>s\geq 1.$
@@ 4.3 Phase 3: Proving $f_{-0}=f_{+0}$
[…]
@@
Let $\alpha\equiv\alpha(\lambda,N,a,m)$ and $\beta\equiv\beta(\lambda,N,a,m)$ be the following polynomials on $\lambda$ of degree NN
@@
$P(\lambda)$= $\left(\lambda+\left[\frac{a+m-1}{2}\right]\right)\left[\frac{a+m}{2}\right]\alpha(\lambda,N,a,m)\beta(\lambda,N,a,-m)$ (4.24)
@@
The first one is straightforward, since $\alpha$ and $\beta$ have degree $N$ by (4.22) and (4.23), and since $Q$ is the product of $2N+1$ monomials.
@@
$(\Rightarrow)$ Suppose that (4.19) holds. Let us show that both (4.20) and (4.21) hold. We divide this direction of the proof into the following three cases:
@@
[…]
@@
[…]
@@
$\times{}_{3}F_{2}\left(\begin{matrix}-q,m-N,-N\\ -\left[\frac{a+m+2}{2}\right]+m-q+1\end{matrix},-2N;1\right).$
@@
(2) Now, suppose $m+N\leq a\leq m+2N+1$. If (4.19) is satisfied, then clearly $\delta_{a-m+2}=0$. Since the left-hand side of (4.19) has degree ama-m, this term comes only from the right-hand side. A straightforward computation yields to
@@
$\times\sum_{r=0}^{N}\begin{pmatrix}N\\ r\end{pmatrix}\frac{\Gamma(2N-r+1)}{\Gamma(N+1)}\frac{\Gamma(\lambda+a-N+r-1)}{\Gamma(\lambda+a-N-1)}.$
@@
$(N+1)_{N}\sum_{r=0}^{N}\frac{(-N)_{r}(\lambda+a-N-1)_{r}}{(-2N)_{r}r!}$ $=(N+1)_{N}{}_{2}F_{1}\left(\begin{matrix}-N,\lambda+a-N-1\\ -2N\end{matrix};\;1\right)$
@@
$1\leq 1+\left[\frac{N}{2}\right]\leq\left[\frac{a-m+2}{2}\right]\leq\left[\frac{2N+3}{2}\right]=N+1.$
@@
Note that if $0<\eta<1$, $\beta_{\eta}(\lambda,N,a,\pm m)$ can be expressed by using the function $_{3}F_{2}$ as:
@@
$(\Leftarrow)$ Let us show now that identity (4.19) is satisfied assuming that (4.20) and (4.21) hold. We do it in two steps. First, we make use of Lemma 8.10 and rewrite the right-hand side of (4.19). Then, we simplify both sides by using properties of the hypergeometric function $_{3}F_{2}$ to show that they actually coincide. From (4.21), we have that (4.19) amounts to:
@@
For $\lambda=N+1-a-q\in\Lambda(N,a)$ and $r=0,\ldots,N$, we set $b:=-\lambda-N+1+r=a+r+q-2N$ and $\ell:=a+m-2N+2r$. From $a\geq m$ and $q\geq\max(0,N-a+m)$, we clearly have $b,\ell\in{\mathbb{N}}$. Now, if $2b-\ell=a-m-2N+2q\geq 0$, then the right-hand side of (4.34) amounts to (4.35) due to Lemma 8.10. On the other hand, if $2b-\ell=a-m-2N+2q<0$, then (4.35) vanishes since all Gegenbauer polynomials in the sum have negative degrees. Similarly, the right-hand side of (4.34) is also zero. In fact, as $m\leq a<m+2N-2q$, we have
@@
We start by showing (4.37). By changing the order of the indices in the sum $S_{1}$, we have
@@
${}_{3}F_{2}\left(\begin{matrix}-d,-N,m-N\\ -2N,1-d+m\end{matrix};\;1\right)=\frac{(-N)_{d}(-m-N)_{d}}{(-m)_{d}(-2N)_{d}},$
@@
If $0\leq q\leq N-1$, (4.38) is trivially satisfied since both sides are zero. Suppose then $N\leq q\leq 2N$. If $d\geq N+1$, then $d-N-1\geq 0$. Therefore, $(-N)_{d}=(-N)(1-N)\cdots(d-N-1)=0$. In other words, the sum $S_{2}$ only has one term, namely, the one for $d=N$. Thus,